# Algorithms for Optimizing Acyclic Queries

### v2

@@ Abstract
Most research on query optimization has centered on binary join algorithms like hash join and sort-merge join. However, recent years have seen growing interest in theoretically optimal algorithms, notably Yannakakis’ algorithm. These algorithms rely on join trees, which differ from the operator trees for binary joins and require new optimization techniques. We propose three approaches to constructing join trees for acyclic queries. First, we give an algorithm to enumerate all join trees of an $\alpha$-acyclic query by edits with amortized constant delay, which forms the basis of a cost-based optimizer for acyclic joins. Second, we show that the Maximum Cardinality Search algorithm by Tarjan and Yannakakis constructs a unique shallowest join tree, rooted at any relation, for a Berge-acyclic query; this tree enables parallel execution of large join queries. Finally, we prove that any connected left-deep linear plan for a $\gamma$-acyclic query can be converted into a join tree by a simple algorithm, allowing reuse of optimization infrastructure developed for binary joins.
@@ 1 Introduction
The query optimizer sits at the heart of a database system. It takes a query as input and generates a plan for efficient execution, allowing users to program declaratively without worrying about evaluation. Among the many relational algebra operators, join has received significant attention in optimization research. Its compositional nature allows for combining information from multiple relations, constructing complex queries from simple ones, and producing an output asymptotically larger than the inputs. The primary challenge is the join ordering problem to find the best arrangement of many join operations. Most existing research has focused on binary join algorithms such as hash join and sort-merge join, but these can produce unnecessarily large intermediates. Recent work has revived interest in optimal join algorithms, notably Yannakakis’ instance-optimal algorithm [44] for acyclic queries, which runs in linear time in the input and output size, $\mathcal{O}(|\textsf{IN}|+|\textsf{OUT}|)$. Its execution is guided by join trees whose nodes are relations, different from traditional binary join plans with relations at the leaves and join operators at the internal nodes. Although the algorithm is optimal regardless of the join tree, the choice of plan can affect practical performance. In this paper, we study the optimization problem in the context of Yannakakis-style algorithms.
@@
  • An algorithm that enumerates all join trees of an acyclic query by edits, with amortized constant time delay.
  • A proof that Maximum Cardinality Search [40] yields a unique shallowest join tree for any Berge-acyclic query, enabling parallel execution for very large queries.
  • A characterization of when the algorithm by Hu et al. [20] can convert a binary join plan into a valid join tree, allowing reuse of existing optimizers.
The rest of the paper is organized as follows: Section˜2 discusses related work; Section˜3 introduces relevant concepts and notations; Section˜4 presents the join tree enumeration algorithm; Section˜5 introduces canonical join trees and their construction; Section˜6 discusses the conversion from binary plans to join trees; finally, Section˜7 points to avenues for future work and concludes the paper. All missing proofs can be found in the appendix, which also contains additional figures, tables and algorithms.
@@ 2 Related Work
Join order optimization is well studied, with algorithms based on dynamic programming (DP) from the bottom up [28, 34, 30], cost-based pruning from the top down [11, 15], greedy heuristics [5, 14, 38], and randomized search [36]. Since the plan space is exponential, most methods prune it: some restrict to left-deep plans [21, 23], while others avoid Cartesian products [28, 29]. Our algorithms restrict the query plans to those running in linear time for acyclic queries. In particular, while avoiding Cartesian products requires each subplan to form a spanning tree of the corresponding subquery’s join graph, [^1] The join graph of a query has a vertex for each relation and an edge for each pair of relations that join with each other. We later define this as the line graph of the query hypergraph in Section 3. our algorithms find maximum spanning trees of the weighted join graph. Several algorithms for ordering binary joins are based on dynamic programming and tabulate shared structures among different plans [28, 29]. This is desirable because subplans are grouped into equivalence classes, and the optimal plan can be constructed in a bottom-up manner. Our enumeration algorithm avoids redundant work by enumerating join trees by edits, i.e., it outputs the difference between consecutive join trees instead of the trees themselves. Future work may explore constructing compact representations of join trees, suitable for dynamic programming.
@@ 3 Preliminaries
We focus on full conjunctive queries [1] in this paper and identify each query with its hypergraph, where there is a vertex for each variable and a hyperedge for each relation.
@@
A hypergraph is a tuple $(X,R,\chi)$ where $X$ is a set of vertices, $R$ a set of hyperedges, and $\chi:R\to 2^{X}$ is a function that assigns to each hyperedge a set of vertices.
Without loss of generality, we limit the scope of our discussion to queries which admit connected hypergraphs. For queries with disconnected hypergraphs, we treat each component separately. We do not consider isolated vertices, empty hyperedges or duplicated hyperedges containing the same set of vertices. Therefore, each hyperedge can be identified with its vertex set. We will use $x\in r$ interchangeably with $x\in\chi(r)$ and apply the common set operations to hyperedges. The size of a hypergraph is the sum of all hyperedges $|H|=\sum_{r\in R}|r|$. We let $H|_{x}$ denote the neighborhood of $x$ in $H$, consisting of all hyperedges containing vertex $x$.
Input: Hypergraph $H$, root $r$
@@
[…]
4while $R\neq\emptyset$ do
5 $i\leftarrow i+1$
6 $r_{i}\leftarrow r$
7 for $x\in r_{i}\cap X$ do
@@ 6 $r_{i}\leftarrow r$
9       $p(r^{\prime})\leftarrow r_{i}$
10
11
12 $X\leftarrow X\setminus r_{i}$
13 $R\leftarrow R\setminus\{r_{i}\}$
@@
Definition 2.
The line graph $L(H)$ of a hypergraph $H=(X,R,\chi)$ is an undirected simple graph $(R,E)$ whose vertices $R$ are the hyperedges of $H$, and there is an edge $(r_{1},r_{2})\in E$ whenever $r_{1}\cap r_{2}\neq\emptyset$. We extend $\chi$ such that $\chi(e)=\chi(r_{1},r_{2}):=r_{1}\cap r_{2}$ for each edge $e=(r_{1},r_{2})\in E$, and define the edge-weight function $\omega(r_{1},r_{2}):=|r_{1}\cap r_{2}|$.
When there is no ambiguity from the context, we simply write $L(H)$ as $L$. The size of a line graph is the sum of all edge weights $|L|=\sum_{e\in E(L)}\omega(e)$. It reflects the number of join clauses in a SQL query with one clause per variable shared between two relations. We let $L|_{x}$ denote the subgraph of $L(H)$ induced by $H|_{x}$.
@@
Figure˜2(a) shows an example hypergraph of size 14. The line graph $L_{6}=L(H_{6})$ is shown in Figure˜2(b). For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph. The edge weight is given by $\omega(Y,U)=|\{a,d\}|=2$. Each unannotated edge in $L_{6}$ connects two hyperedges sharing only the vertex $a$. Otherwise, the common vertices shared by a pair of hyperedges are annotated next to the corresponding edge. The size of the line graph is $|L_{6}|=19$.
Definition 4.
A join tree $T$ of hypergraph $H$ is a spanning tree of $L(H)$ such that $T|_{x}$ is a connected subtree for each $x\in X(H)$. If a certain vertex is specified as the root, $T$ becomes a rooted join tree where the edges are oriented away from the root.
A join tree $T(L)$ is a subgraph of $L$, so we use $R(T)$ to denote the set of tree nodes and $E(T)$ to denote the set of tree edges. Because $T$ is spanning and acyclic, we always have $R(T)\equiv R(L)$ and $E(T)\subseteq E(L)$. The requirement that $T|_{x\in X(H)}$ is connected is also known as the running intersection property [13]. We write $\mathcal{T}(L(H))$ or $\mathcal{T}(H)$ to denote the set of unrooted join trees. The union join graph $\mathcal{U}(L(H))$ or $\mathcal{U}(H)$ is a spanning subgraph of $L(H)$ where the edge set is the union of all join trees $E(\mathcal{U}(H))=\bigcup_{T\in\mathcal{T}(H)}E(T)$. We use $T_{r}$ to denote a join tree rooted at $r\in R(T)$. When there is no ambiguity, we simply write $T$. The depth of a node $r_{i}$ in the rooted tree $T_{r}$, denoted $0pt(T_{r},r_{i})$, is defined as its distance from the root. Join trees can be constructed by a procedure called GYO reduction.
Definition 5.
The GYO reduction algorithm finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree. It generates the join tree $T^{G}$ as shown in Figure˜2(c). Readers may refer to the work of Yu and Ozsoyoglu [45] for details.
Four common notions of hypergraph acyclicity are defined in decreasing order of strictness, following the seminal work by Fagin [13], namely Berge-acyclic $\Rightarrow$ $\gamma$-acyclic $\Rightarrow$ $\beta$-acyclic $\Rightarrow$ $\alpha$-acyclic. The detailed definitions are given by Appendix˜A in Appendix˜A.
Another algorithm for checking acyclicity of hypergraphs and constructing join trees is the Maximum Cardinality Search (MCS) algorithm by Tarjan and Yannakakis [40]. Algorithm˜1 shows a simplified pseudocode of MCS. We discuss the properties of MCS and how it constructs the join trees $T^{M}$ in Figure˜2(d) with Appendix˜A in Appendix˜A.

### v3

@@ Abstract
Most research on query optimization has centered on binary join algorithms like hash join and sort-merge join. However, recent years have seen growing interest in theoretically optimal algorithms, notably Yannakakis’ algorithm. These algorithms rely on join trees, which differ from the operator trees for binary joins and require new optimization techniques. We propose three approaches to constructing join trees for acyclic queries. First, we give an algorithm to enumerate all join trees of an $\alpha$-acyclic query by edits in linear time with amortized constant delay, which forms the basis of a cost-based optimizer for acyclic joins. Second, we show the Maximum Cardinality Search algorithm by Tarjan and Yannakakis constructs the unique shallowest join tree for any Berge-acyclic query, thus enabling parallel execution of large join queries. Finally, we prove that a simple algorithm by Hu et al. converts any connected left-deep linear plan of a $\gamma$-acyclic query into a join tree, allowing reuse of optimizers developed for binary joins.
@@ 1 Introduction
The query optimizer sits at the heart of a database system. It takes a query as input and generates a plan for efficient execution, allowing the user to program declaratively without worrying about performance. Among the many relational algebra operators, join has received significant attention in optimization research. Its compositional nature allows for combining information from multiple relations, constructing complex queries from simple ones, and producing an output asymptotically larger than the inputs. The primary challenge is the join ordering problem to find the best arrangement of many join operations. Most existing research has focused on binary join algorithms such as hash join and sort-merge join, but these can produce unnecessarily large intermediates. Recent work has revived interest in optimal join algorithms, notably Yannakakis’ instance-optimal algorithm [45] for acyclic queries, which runs in linear time in the input and output size, $\mathcal{O}(|\textsf{IN}|+|\textsf{OUT}|)$. Its execution is guided by join trees whose nodes are relations, different from traditional binary join plans with relations at the leaves and join operators at the internal nodes. Although the algorithm is optimal regardless of the join tree, the choice of plan can affect practical performance. In this paper, we study the optimization problem in the context of Yannakakis-style algorithms.
@@
  1. 1.
    Given the line graph $L$ (section˜3) of an $\alpha$-acyclic query $H$, algorithm˜2 enumerates all join trees by edits [^1] To avoid redundant work, enumeration by edits [24] outputs the difference between consecutive elements. with amortized constant delay, i.e., in $\mathcal{O}(|L|+k)$ where $k$ is the number of trees generated (Theorem˜27). If the query is $\gamma$-acyclic, algorithm˜4 further reduces the total time complexity to $\mathcal{O}(|H|+k)$ (Theorem˜28).
  2. 2.
    Given a Berge-acyclic query, we prove that the classic Maximum Cardinality Search algorithm by Tarjan and Yannakakis [41] yields a unique shallowest join tree (Theorem˜35), enabling parallel execution of very large queries.
  3. 3.
    Given a connected left-deep linear join plan for any $\gamma$-acyclic query, we prove that a simple algorithm by Hu et al. [22] always converts the plan into a valid join tree (section˜6), allowing reuse of existing optimizers.
The rest of the paper is organized as follows: section˜2 discusses related work; section˜3 introduces relevant concepts and notations; section˜4 presents the join tree enumeration algorithms; section˜5 introduces the unique shallowest join tree, namely the canonical join tree and its construction; section˜6 discusses the conversion from binary plans to join trees; section˜7 concludes the paper by pointing to avenues for future work. Additional technical details, proofs and a notation table (table˜2) can be found in the appendix.
@@ 2 Related Work
Join order optimization is well studied, with algorithms based on dynamic programming (DP) from the bottom up [30, 36, 32], cost-based rewriting from the top down [13, 17], greedy heuristics [6, 16, 40], and randomized search [38]. Since the plan space is exponential, most methods prune it: some restrict to left-deep plans [23, 25], while others avoid Cartesian products [30, 31]. Our algorithms restrict the query plans to those running in linear time for acyclic queries. In particular, while avoiding Cartesian products requires each subplan to form a spanning tree of the corresponding subquery’s join graph, [^2] The join graph of a query has a vertex for each relation and an edge for each pair of relations that join with each other. We later define this as the line graph of the query hypergraph in section 3. our algorithms find maximum spanning trees of the weighted join graph. Several algorithms for ordering binary joins are based on dynamic programming and tabulate shared structures among different plans [30, 31]. This is desirable because subplans are grouped into equivalence classes, and the optimal plan can be constructed in a bottom-up manner. Future work may explore constructing compact representations of join trees, suitable for dynamic programming.
@@ 3 Preliminaries
We focus on full conjunctive queries [1] in this paper and identify each query with its hypergraph, where each vertex represents a variable and each hyperedge represents a relation.
Definition 1 (Hypergraph).
A hypergraph $H=(X,R,\chi)$ consists of a set of vertices $X$, a set of hyperedges $R$ and an incidence function $\chi:R\to\mathcal{P}(X)$.
We only consider hypergraphs without isolated vertices, empty hyperedges or duplicated hyperedges over the same set of vertices. We assume each hyperedge $r$ contains a bounded number ($\mathcal{O}(1)$) of vertices exclusive to $r$. When there is no ambiguity, we will use $r$ interchangeably with $\chi(r)$, $x\in r$ with $x\in\chi(r)$, and $H(X,R)$ with $H(X,R,\chi)$. We will also apply common set operations directly to hyperedges, e.g., $r_{1}\cap r_{2}$ for $\chi(r_{1})\cap\chi(r_{2})$. In fact, the reader can often ignore $\chi$ and identify a hyperedge with its set of vertices; the purpose of $\chi$ is to relate hyperedges over different vertex sets across hypergraphs in the correctness proofs of our algorithms. We write $X(H)$ and $R(H)$ to refer to the vertices and edges of $H$ respectively. The size of a hypergraph is the total size of all hyperedges $|H|=\sum_{r\in R}|r|$. We let $H|_{x}$ denote the neighborhood of $x\in X$, consisting of all hyperedges containing $x$.
Definition 2 (Multigraph).
A multigraph $G=(R,E,\rho)$ consists of a set of vertices $R$, a set of edges $E$, and the incidence function $\rho:E\to\mathcal{P}(R)$ such that $1\leq|\rho(e)|\leq 2$ for any $e\in E$. An edge $e$ is a self-loop if $|\rho(e)|=1$. Edges $e_{1},e_{2}\in E$ are parallel if $\rho(e_{1})=\rho(e_{2})$.
We denote the set of vertices in $G$ with $R$, because we will soon define the line graph $L$ of a hypergraph where each vertex in $L$ represents a hyperedge. We write $R(G)$ and $E(G)$ to refer to the vertices and edges of $G$ respectively.
Definition 3 (Simple Graph).
A simple graph is a multigraph with no parallel edges or self-loops, i.e., the incidence function is injective and always returns two distinct vertices.
Because each edge $e$ in a simple graph can be identified with its two endpoints $\rho(e)=\{r_{1},r_{2}\}$, we will omit $\rho$ and write $\{r_{1},r_{2}\}$ for $e$ when there is no ambiguity.
Definition 4 (Cycle, Clique and Diamond).
In a simple graph $G=(R,E)$, a sequence of $n\geq 3$ distinct vertices $(v_{0},\ldots,v_{n-1})$ where $v_{i}\in R$ is
  • a cycle if there is an edge between $v_{i}$ and $v_{(i+1)\bmod n}$ for all $0\leq i\leq n-1$;
  • an n-clique $K_{n}$ if the induced graph $G|_{\{v_{0},\dots,v_{n}\}}$ has an edge between each pair of vertices;
  • a diamond if $n=4$ and $G|_{\{v_{0},\dots,v_{3}\}}$ has one fewer edge than a $K_{4}$.
A weighted graph is a (multi-)graph where each edge is assigned a weight:
Definition 5 (Weighted Graph).
A weighted graph $(G,\omega)$ consists of a multigraph $G=(R,E,\rho)$ and weight function $\omega:E\to\mathbb{N}$ assigning a natural number to each edge in $G$.
The line graph of a hypergraph is the “intersection graph” of its hyperedges:
Definition 6 (Line Graph).
For a hypergraph $H$, the line graph $L(H)=(G,\chi)$ consists of a simple graph $G=(R,E)$ and an edge labeling function $\chi:E\to\mathcal{P}(X)$. The vertices of $G$ are the hyperedges $R$ of $H$, and there is an edge $e=\{r_{1},r_{2}\}\in E$ when $r_{1}\cap r_{2}\neq\emptyset$. The edge labeling function maps each edge $e$ to the intersection of the hyperedges represented by the endpoints of $e$: $\chi(\{r_{1},r_{2}\})=r_{1}\cap r_{2}$.
In addition, we define a weight function mapping each edge to the size of its label:
Definition 7 (Line Graph Edge Weight).
For each edge $e\in E(L(H))$, the weight function $\omega:E(L(H))\to\mathbb{N}$ is defined as $\omega(e)=|\chi(e)|$.
When there is no ambiguity from the context, we simply write $L(H)$ as $L$, and we may treat the line graph $(G,\chi)$ as a weighted graph $(G,\omega)$. The size of a line graph is the sum of all edge weights $|L|=\sum_{e\in E(L)}\omega(e)$. For any subgraph $G$ of $L(H)$, we let $G|_{x}$ denote the subgraph of $G$ induced by $H|_{x}$.
Example 8.
fig.˜2(a) shows an example hypergraph of size 14. The line graph $L_{6}=L(H_{6})$ is shown in fig.˜2(b). For instance, hyperedges $Y$ and $U$ share two vertices $a$ and $d$, so they are connected by an edge in the line graph. The edge weight is given by $\omega(\{Y,U\})=|\{a,d\}|=2$. Each unannotated edge in $L_{6}$ connects two hyperedges sharing only the vertex $a$. Otherwise, the common vertices shared by a pair of hyperedges are annotated next to the corresponding edge. The size of the line graph is $|L_{6}|=19$.
Definition 9 (Join Tree).
A join tree $T$ of hypergraph $H$ is a spanning tree of $L(H)$ such that $T|_{x}$ is a connected subtree for each $x\in X(H)$. If a certain vertex is specified as the root, $T$ becomes a rooted join tree.
We use $R(T)$ to denote the set of nodes and $E(T)$ to denote the set of edges of tree $T$. The requirement that $T|_{x\in X(H)}$ be a connected subtree is also known as the running intersection property [15]. We write $\mathcal{T}(L(H))$ or $\mathcal{T}(H)$ to denote the set of unrooted join trees of $H$. We use $T_{r}$ to denote a join tree rooted at $r\in R(T)$. When there is no ambiguity, we simply write $T$. The depth of a node $r_{i}$ in the rooted tree $T_{r}$, denoted $0pt(T_{r},r_{i})$, is defined as its distance from the root.
Four common notions of hypergraph acyclicity are defined in increasing order of strictness [15], namely $\alpha$-acyclic $\supset$ $\beta$-acyclic $\supset$ $\gamma$-acyclic $\supset$ Berge-acyclic.
Definition 10 (Hypergraph Acyclicity).
A hypergraph $H$ is:
  • $\alpha$-acyclic if it admits a join tree as in section˜3;
  • $\beta$-acyclic if every subgraph of $H$ is $\alpha$-acyclic;
The following classic result relates join trees of $H$ to maximum spanning trees of $L(H)$:
Theorem 11 (Maier [29]).
Given an $\alpha$-acyclic hypergraph $H=(X,R)$, a tree with nodes in $R$ is a join tree of $H$ if and only if it is a maximum spanning tree of $L(H)$.
Join trees for acyclic hypergraphs can be constructed by a procedure called GYO reduction.
@@
The GYO reduction algorithm [46] finds such an order iteratively, and attaches each hyperedge to its parent to form a join tree. It generates the join tree $T^{G}$ as shown in fig.˜2(c), where, for example, $T$ is the parent of $W$, denoted as $p(W)=T$ and the root $P$ has no parent.
Input: $\alpha$-acyclic hypergraph $H$ and root $r$
@@
3while $R\neq\emptyset$ do
4
5 for $x\in r\cap X$ do
@@
7       $p(r^{\prime})\leftarrow r$
8
9
10 $X\leftarrow X\setminus r$
11 $R\leftarrow R\setminus\{r\}$
12 $E(T_{r})\leftarrow E(T_{r})\cup\{\{r,p(r)\}\}$
@@
Another way to construct join trees is via the Maximum Cardinality Search (MCS) algorithm [41], and we present a simplified version in algorithm˜1. The join tree produced by the algorithm is called an MCS tree. Given an input hypergraph $H$ and a hyperedge $r\in R(H)$ chosen as the root, algorithm˜1 constructs a rooted MCS tree by iteratively expanding toward hyperedges connected with those already processed. The algorithm maintains a set $R$ of unprocessed hyperedges and a set $X$ of unvisited vertices. Initially, the MCS tree contains no edges, and the first hyperedge to be processed is the root $r$. In each iteration, the algorithm processes the current hyperedge $r$ by examining all of its unvisited vertices in $r\cap X$. Whenever such a vertex $x$ is encountered, every remaining hyperedge $r^{\prime}\in R$ that contains $x$ records $r$ as its tentative parent $p(r^{\prime})\leftarrow r$. Once all vertices of $r$ are marked as visited, they are removed from $X$. The hyperedge $r$ is marked as processed and removed from $R$. After processing each $r$, the edge between $r$ and its recorded parent $p(r)$ is added to the MCS tree as $\{r,p(r)\}$. The algorithm selects as the next hyperedge an element of $R$ that maximizes the number of already visited vertices, namely a hyperedge $r$ that maximizes $|r\setminus X|$, breaking ties arbitrarily. This process continues until all hyperedges have been processed, when the recorded edges $\{r,p(r)\}$ form an MCS tree rooted at $r$.
@@
On $H_{6}$ in fig.˜2(a) with $r=P$, algorithm˜1 examines all vertices of $P$. $P$ also becomes the tentative parent of five hyperedges, $p(S)=p(T)=p(U)=p(W)=p(Y)\leftarrow P$. It proceeds to check the next hyperedge with the most visited vertices as in Line 1. At this point, each of the five has one visited vertex, algorithm˜1 breaks the tie arbitrarily, say proceeding with $S$, and examining all its vertices. Its parent is finalized as $p(S)=P$. Each of the remaining four again has one visited vertex, so algorithm˜1 proceeds to check $T$. Its parent is finalized as $p(T)=P$. While checking $T$, algorithm˜1 examines the vertex $c$ that is also in $U,W$, their parents are updated to $p(U)=p(W)\leftarrow T$. It continues until all hyperedges are processed. By connecting each hyperedge with its parent, algorithm˜1 constructs the MCS tree $T^{M}$ in fig.˜2(d).
algorithm˜1 can generate different join trees by varying the choice of root and tie-breaking strategy. However, it does not guarantee to generate all possible trees. For example, it never generates the join tree $T^{G}$ shown in fig.˜2(c).
We extend the notions of parent, children and siblings to edges of a rooted tree. Let $e\in E(T)$ be an edge of tree $T$ rooted at $r$ (and $e$ not incident to $r$). Its parent $p(e)\in E(T)$ is the unique tree edge incident to $e$ and lying closer to the root $r$. We let $c(e)$ denote the set of children of $e$ in $T$. The siblings of $e$ is the set of edges incident to the parent of $e$, namely $s(e)=\left\{e^{\prime}\in E(T)\setminus\{e\}\mid p(e^{\prime})=p(e)\right\}$. Edges incident to the root do not have a parent and are all siblings.
section˜3 shows that an MCS tree is “somewhat monotonic” in that every edge must contain some variable not in its parent. If two edges share any variable not in their parents, they must be siblings.
Lemma 14.
Let $T$ be an MCS tree. For an edge $e\in T$ that has a parent, then
  1. 1.
    $e\not\subseteq p(e)$
  2. 2.
    For another edge $e^{\prime}\in T:(e\setminus p(e))\ \cap\ (e^{\prime}\setminus p(e^{\prime}))\ \neq\ \emptyset\quad\implies\quad p(e)=p(e^{\prime}).$